# UFORecon: Generalizable Sparse-View Surface Reconstruction from Arbitrary and UnFavOrable Data Sets

### 08 Mar 2024 v1

@@ 3 Motivation
Figure 3: Overall pipeline of UFORecon. Our cross-view matching transformer extracts cross-view matching features from multi-level image features of the given image set (Sec. 4.2). Cross-view matching features are then represented as 3D volumes (i.e., cascaded correlation frustums, (Sec. 4.3) and as the 2D features (i.e., geometry aware similarity encoding, (Sec. 4.4). Our reconstruction transformer (Sec. 4.5) effectively fuses various representations of matching features and geometry features (i.e.,  depth) for volume rendering and color blending.
@@ 5 Experiments
Table 1: A quantitative results on favorable sets. Our method outperforms traditional 3D reconstruction methods (first section), per-scene surface reconstruction methods (second section), generalizable novel-view synthesis methods (third section), and generalizable surface reconstruction methods (fourth section). The Bold numbers mean the best, and the underlined values indicate the second-best scores.
@@
Table 2: A quantitative results on unfavorable sets. The Bold numbers mean best and underlined numbers mean 2nd best scores.
@@ 5.1 Experimental Settings
Baseline. To demonstrate the effectiveness of our method, we mainly compare it with (1) SparseNeus [29], VolRecon  [40], ReTR [25], and C2F2NeuS [54] the state-of-the-art generalizable neural implicit reconstruction method. For evaluating scene and viewpoints generalizability, we additionally compare with (2) Generalizable neural rendering methods [62, 4, 49], and (3) Neural implicit reconstruction [58, 36, 61, 47] which require per-scene training. Further details about the baselines are provided in the Appendix.
@@
Figure 6: Effect of random set training on unfavorable sets. With the random set training, the quality is much smoother and complete.
@@ 6 Conclusion
Limitations and future work. Although our method achieves advancements in view-combination generalizability for surface reconstruction, it is still challenging to reconstruct scenes with much higher complexity. In addition, We hope that our contributions will inspire further research to extend this to such directions.
@@ B Detailed Description of Baseline Methods
In our research, we compared our method with four different types of surface reconstruction approaches across all experiments. (1) The most directly comparable category is generalizable implicit surface reconstruction [30, 39, 26, 55], where our method is included. These approaches assume a very limited number of views and unseen scenes during inference. (2) Generalizable neural rendering methods [63, 50, 5] also use a very limited number of views to perform novel view synthesis from unseen scenes. These methods learn geometry through volume density rather than SDF, leading to ambiguity as they lack an explicit surface definition. By applying thresholding on volume density and with Marching Cubes [31], geometry can be obtained. (3) Neural implicit surface reconstruction [59, 60, 35, 48] techniques extract accurate geometry from densely captured multi-view images of a single scene via optimization. Since these methods cannot generalize across multiple scenes, per-scene optimization is required to estimate geometry, unlike generalizable methods. (4) We compare with a conventional matching-based method, COLMAP [43]. It utilizes feature matching and triangulation from multiple images, providing a baseline for comparing the effectiveness of neural network-based methods.
@@ F Impact of the Number of Views
G More Qualitative Results

### 11 Mar 2024 v2

# UFORecon: Generalizable Sparse-View Surface Reconstruction from Arbitrary and UnFavOrable Sets

@@ 3 Motivation
Figure 3: Overall pipeline of UFORecon. Our cross-view matching transformer extracts cross-view matching features from multi-level image features of the given image set (Sec. 4.2). Cross-view matching features are then represented as 3D volumes (i.e., cascaded correlation frustums, (Sec. 4.3) and as the 2D features (i.e., geometry aware similarity encoding, (Sec. 4.4). Our reconstruction transformer (Sec. 4.5) fuses various representations of matching features and geometry features (i.e.,  depth) for volume rendering and color blending.
@@ 5 Experiments
Table 1: A quantitative results on favorable sets. Our method outperforms traditional 3D reconstruction methods (first section), per-scene surface reconstruction methods (second section), generalizable novel-view synthesis methods (third section), and generalizable surface reconstruction methods (fourth section). The Bold numbers mean the best, and the underlined values indicate the second-best scores. (*) denotes the use of a random set training strategy.
@@
Table 2: A quantitative results on unfavorable sets. (*) denotes the use of a random set training strategy.
@@ 5.1 Experimental Settings
Baseline. To demonstrate the effectiveness of our method, we mainly compare it with (1) SparseNeus [28], VolRecon  [35], ReTR [26], and C2F2NeuS [47] the state-of-the-art generalizable neural implicit reconstruction method. For evaluating scene and viewpoints generalizability, we additionally compare with (2) Generalizable neural rendering methods [51, 4, 43], and (3) Neural implicit reconstruction [50, 42] which require per-scene training. Additionally, we include (4) traditional MVS methods [37, 9].
@@
Figure 6: Effect of random set training on unfavorable sets.
@@ 6 Conclusion
Limitations and future work. Although our method achieves advancements in view-combination generalizability for surface reconstruction, it is still challenging to reconstruct scenes with much higher complexity. In addition, we assume that the ground-truth camera poses are given. We include a detailed discussion on this in the Appendix. We hope that our contributions will inspire further research to extend this to such directions.
@@ B Detailed Description of Baseline Methods
In our research, we compared our method with four different types of surface reconstruction approaches across all experiments. (1) The most directly comparable category is generalizable implicit surface reconstruction [28, 35, 26, 47], where our method is included. These approaches assume a very limited number of views and unseen scenes during inference. (2) Generalizable neural rendering methods [51, 43, 4] also use a very limited number of views to perform novel view synthesis from unseen scenes. These methods learn geometry through volume density rather than SDF, leading to ambiguity as they lack an explicit surface definition. By applying thresholding on volume density and with Marching Cubes [marchingcubes], geometry can be obtained. (3) Neural implicit surface reconstruction [50, 42] techniques extract accurate geometry from densely captured multi-view images of a single scene via optimization. Since these methods cannot generalize across multiple scenes, per-scene optimization is required to estimate geometry, unlike generalizable methods. (4) We compare with a conventional matching-based method, COLMAP [37]. It utilizes feature matching and triangulation from multiple images, providing a baseline for comparing the effectiveness of neural network-based methods. (5) Additionally, we include learning-based MVS methods [48, 9].
@@ F Impact of the Number of Views
G Comparison on Depth Estimation.
MVS methods and other baseline methods experience significant performance degradation in unfavorable settings (Table 8), which are more susceptible to self-occlusion. The correlation frustum is one of our key contributions (not introduced in C2F2NeuS) to model global correlation among input views, and the reconstruction transformer estimates SDFs in a correlation-aware manner. We achieve robust performance in handling arbitrary view combinations.
Favorable Unfavorable
Method
<1mm $\uparrow$ <4mm $\uparrow$ Abs. $\downarrow$ <1mm $\uparrow$ <4mm $\uparrow$ Abs. $\downarrow$
MVSNet [43]
29.95 72.33 13.62 5.82 19.72 72.74
TransMVSNet[8]
38.95 82.91 13.53 6.21 15.60 161.71
VolRecon
43.60 82.53 7.12 3.93 11.43 185.95
ReTR [22]
51.95 85.06 5.41 10.83 19.26 58.02
Ours
54.80 86.69 6.09 51.30 65.14 17.02
Table 8: 3-Views Depth Map Comparison on DTU Datasets.
H Comparative Ablation Study on Depth Supervision
As shown in Table 9, our method works well even without depth supervision. Especially in unfavorable sets, our method without depth supervision outperforms the baselines even with depth supervision. Depth supervision in SparseNeuS tends to result in over-smooth surfaces, which is also reported in [35].
[…]
Method
w/o $\mathcal{L}_{d}$ w/ $\mathcal{L}_{d}$ w/o $\mathcal{L}_{d}$ w/ $\mathcal{L}_{d}$
SparseNeuS [24]
1.64 4.22 4.16 5.55
VolRecon [30]
2.06 1.38 8.65 3.18
ReTR [22]
1.45 1.17 4.41 2.94
Ours
1.26 0.99 2.26 1.56
Table 9: Ablation study of Depth Supervision ($\mathcal{L}_{d}$).
I Regarding Pracitcal Relevance.
Apparently, obtaining accurate camera poses from a sparse set of images presents challenges. In response, a parallel line of research has focused on estimating relative camera poses from sparse views [38, 6], and it is anticipated that we can apply these methods to obtain more accurate camera poses from sparse, unfavorable set of views. Another research direction explores reconstructing 3D in a pose-free manner only using source images [10, 17], in which case our proposed modeling of correlation across images becomes even more meaningful. These approaches potentially address concerns on obtaining camera poses. We believe that obtaining accurate geometry under pose-free settings remains underdeveloped and seems to be a promising future work.
J More Qualitative Results